OpenStudy (anonymous):

Which statement shows how two polynomials 5x − 6 and 6x + 2 demonstrate the closure property when multiplied? 30x2 − 26x − 12 may or may not be a polynomial 30x2 − 11x − 12 may or may not be a polynomial 30x2 − 26x − 12 is a polynomial 30x2 − 11x − 12 is a polynomial

OpenStudy (freckles):

First multiply

OpenStudy (anonymous):

wait can u give me an example im not quite sure how to multiply

OpenStudy (anonymous):

it

myininaya (myininaya):

so if i asked you to multiply 2x+1 and 4x-5 you would do the following: by the way 2x+1 and 4x-5 and 5x-6 and 6x+2 are all polynomials by definition of a polynomial so anyways multiplying 2x+1 and 4x-5 \[(2x+1)(4x-5)=2x(4x-5)+1(4x-5)\\ =2x(4x)-2x(5)+1(4x)-1(5) \\ =8x^2-10x+4x-5 \\ =8x^2-6x-5\]

myininaya (myininaya):

is 8x^2-6x-5 a polynomial?

OpenStudy (anonymous):

no

myininaya (myininaya):

why would you say that

myininaya (myininaya):

do you know the definition of a polynomial?

myininaya (myininaya):

\[a_nx^n+a_{n-1}x^{n-1}+...+a_3x^3+a_2x^2+a_1x^1+a_0\] where n is positive integer or zero those ai (the coefficients can be zero) the exponents all have to be positive integers or zero

myininaya (myininaya):

examples of polynomials: 1 5 x x+1 x^2+2x+1 x^3-x-4 x^5+1 x^7 examples of an expression that is not polynomial: x^-1 x^-1+2 sqrt(x) |x| x^2/3+1 1/x -1/(x-1) I could sit here and give you tons of examples of each but you should get the idea of what is a polynomial now and what is not